OpenStudy (anonymous):

How would I graph y = 5/3x + 2?

OpenStudy (callisto):

From the equation, you see y-int =2 So you'll get a point at (0,2) Then put y= 0 into the equation and solve x So, you'll get another point After that, draw a coordinate plane with marked scale Plot the 2 points you got Draw a line joining the 2 points You'll get the graph you want!

OpenStudy (anonymous):

so like this callisto?  would I go even farther out to the four would I go back to -2 or go forward to 2?

OpenStudy (callisto):

Hmm.. obviously not, the graph should have a positive slope but yours has a negative one...

OpenStudy (callisto):

y = 5/3x + 2 Put y=0 into the equation 0 = (5/3)x +2 -2 = (5/3) x x = -6/5